OpenStudy (anonymous):

Find the equation of the tangent line to the graph of f(x)= ln(1+x^2) at the point (0,0)

OpenStudy (zepp):

Can you find the derivative of this function?

OpenStudy (anonymous):

Find derivative and then put it in the eqn: (y-y1) = m(x-x1)

OpenStudy (anonymous):

where m = derivative x = 0 y = 0

OpenStudy (anonymous):

\[f'(0)=m \\ (x_a,y_b)=(0,0) \\ y-y_a=m(x-x_a)\]

OpenStudy (anonymous):

The only real work that it tough is finding the derivative of \[f(x)= \ln(1+x^2) \]Think you can find it?

OpenStudy (zepp):

\[\large \frac{dy}{dx}f(x)=\ln(1+x^2)=\frac{\frac{dy}{dx}(1+x^2)}{1+x^2}=\frac{2x}{1+x^2}\]

OpenStudy (anonymous):

ok the derivative of this function is (derivatives gives us slope) \[m=f'(x)=\frac{2x}{1+x^2}\] put x= 0 in the above to get numerical value of slope \[m=\frac{2*0}{1+0^2}=0\] so slope is zero! and you may know that the slope of horizontal line is zero whose equation is !

OpenStudy (anonymous):

did u know the equation of horizontal line??

OpenStudy (anonymous):

and the asker is just looking around :(

OpenStudy (anonymous):

@cunninnc

OpenStudy (anonymous):

@sami-21 yes thank you

OpenStudy (anonymous):

to find the tangent line of this eq. \[f(x)=e ^{x}/2x+1 \] would my derivative be \[f \prime= (2x+1)(e ^{x}) - (e ^{x}) (2) / (2x+1)^{2} \]

OpenStudy (anonymous):

\[f′=\frac{(2x+1)(e^x)−(e^x)(2)}{(2x+1)^2}\]

OpenStudy (anonymous):

Yup you are right

OpenStudy (anonymous):

thanks so do i leave it like that or simplify

OpenStudy (anonymous):

Do i simply solve by pluggin in?